A toboggan approaches a snowy hill moving at 10.9 m/s. The coefficients of static and kinetic friction between the snow and the toboggan are 0.50 and 0.35, respectively, and the hill slopes upward at 40.0 ∘ above the horizontal.

A) Find the magnitude of the acceleration of the toboggan as it is going up the hill.

B) What is the direction of the acceleration as the toboggan is going up the hill?

C) Find the magnitude of the acceleration of the toboggan after it has reached its highest point and is sliding down the hill.

D) What is the direction of the acceleration as the toboggan is sliding down the hill?
Answer #1

A)

Acceleration, a = g(sinθ + μk cosθ)

a = 9.8(sin40 + 0.35 cos40) = 8.93 m/s^2

B)

Up the plane

C)

Acceleration, a = g(sinθ - μk cosθ)

a = 9.8(sin40 - 0.35 cos40) = 3.67 m/s^2

D)

Down the plane

Comment in case any doubt please rate my answer...